### v1

@@ 7 Conclusion
This paper presented a unified framework for designing ephemeris and almanac messages for lunar navigation satellites, addressing the unique dynamical and operational challenges of the lunar environment. By combining osculating orbital elements with Chebyshev and Fourier representations, the proposed ephemeris model achieves high-fidelity orbit reconstruction over extended fitting intervals while satisfying the LunaNet message size constraint. Across multiple orbit regimes—including 30-, 24-, and 12-hour ELFOs and a 6-hour polar orbit—the hybrid representation achieved sub-meter position and sub-millimeter-per-second velocity fitting accuracy within 900 bits for ephemeris arcs up to four hours.
For the almanac, a compact polynomial–Fourier representation of the orbital elements successfully captured long-term orbital variations over 15-day intervals, yielding median latitude and longitude errors below one degree. This level of accuracy is sufficient for distinguishing visible satellites during warm-start operations, ensuring robust acquisition performance after lunar night periods.
Overall, the proposed approach provides a scalable and interoperable foundation for generating almanac and ephemeris products for the Lunar Augmented Navigation Service (LANS). Future work will extend this framework to incorporate orbit determination and prediction errors, satellite clock models, and user receiver simulations to evaluate end-to-end navigation performance within the broader LunaNet architecture.

### v2

@@ 7 Conclusion
This paper presented a unified framework for the design of ephemeris and almanac messages for lunar navigation satellites, addressing the unique dynamical and operational challenges of the lunar environment. By combining osculating orbital elements with Chebyshev and Fourier representations, the proposed ephemeris model achieves high-fidelity orbit reconstruction over extended fitting intervals while satisfying the LunaNet message size constraints. Across multiple orbit regimes—including 30-, 24-, and 12-hour ELFOs, as well as a 6-hour polar orbit—the hybrid representation achieves sub-meter position accuracy and sub-millimeter-per-second velocity accuracy within a 900-bit ephemeris budget for fitting intervals of up to four hours.
For the almanac, a compact polynomial–Fourier representation of the orbital elements successfully captures long-term orbital variations over 15-day intervals. This approach yields median elevation and azimuth estimation errors below $2^{\circ}$, as well as range and Doppler shift estimation errors below 10% of their true values, thereby effectively constraining the acquisition search space for lunar navigation signals.
Overall, the proposed framework provides a scalable and interoperable foundation for generating ephemeris and almanac products for LANS. Future work will extend this approach to incorporate orbit determination and prediction uncertainties, satellite clock models, and user receiver simulations in order to evaluate end-to-end navigation performance within the broader LunaNet architecture.
Acknowledgment
AI tools including ChatGPT 5 was used to improve wording and correct grammar/style. AI tools were not used to generate ideas, results, figures, or analyses.